OpenStudy (anonymous):

if f(x)=x^2+4x, find the values of x if the value of f(x) is 5.

OpenStudy (anonymous):

hmmm. i tried that but the answer is (1,-5).

OpenStudy (anonymous):

If F9X) = 5, then x^2 + 4x = 5 So, x^2 + 4x - 5 = 0 (x+5)(x-1) = 0 Therefore, x = -5, 1

jigglypuff314 (jigglypuff314):

easyaspi314 is correct, I had read it wrong the first time

OpenStudy (anonymous):

How did you get F9x in the beginning?

jigglypuff314 (jigglypuff314):

typo they meant f(x)